Charles Lackey operates a bakery in Idaho Falls, Idaho. Because of its excellent product and excellent location, demand has increased by

45 %

in the last year. On far too many occasions, customers have not been able to purchase the bread of their choice. Because of the size of the store, no new ovens can be added. At a staff meeting, one employee suggested ways to load the ovens differently so that more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by hand, requiring additional manpower. This is the only thing to be changed. The bakery currently makes 1500 loaves per month. The pay will be

$8 per hour for employees and each employee works 160 hours per month. Charles Lackey can also improve the yield by purchasing a new blender. The new blender will mean an increase in his investment. This new blender will mean an increase in his costs of $100 per month, but he will achieve the same new output (an increase to

175.00 as the change in labor hours.

a) Current productivity for 640 work hours =

nothing

loaves/dollar (round your response to three decimal places).

If Charles chooses to increase the number of work hours to

928

in order to employ the new oven loading technique, then the productivity is =nothing

loaves/dollar (round your response to three decimal places).

b) If Charles instead chooses to purchase a new blender (while holding labor constant at 640 hours at

$88

per

hour) ,

then the productivity is =

nothing

loaves/dollar (round your response to three decimal places).

c) By adding manpower, the percentage increase in productivity is

nothing %

(enter your response as a percentage rounded to two decimal places and include a minus sign if necessary).

By purchasing a new blender (while holding labor constant at 640 hours at$8 perhour) ,

the percentage increase in productivity is

nothing %

(enter your response as a percentage rounded to two decimal places and include a minus sign if necessary
Answer #1

1. VALUES
LABOR = 5120(640 * 8)
OUTPUT = 1500
PRODUCTIVITY = OUTPUT / TOTAL INPUT
PRODUCTIVITY = 1500 / (5120) = 0.293

B. VALUES
LABOR = 7424(728 * 8)
OUTPUT = 2175(45% INCREASE)
PRODUCTIVITY = OUTPUT / TOTAL INPUT
PRODUCTIVITY = 2175 / (7424) = 0.293

2. VALUES
LABOR = 5120
OVERHEAD = 100
OUTPUT = 2175
PRODUCTIVITY = OUTPUT / TOTAL INPUT
PRODUCTIVITY = 2175 / (5120 + 100) = 0.417
.
3. A. SINCE PRODUCTIVITY IN BOTH THE CASES IS THE SAME, THERE WILL BE NO INCREASE OR DECREASE.

B. CHANGE IN PERCENTAGE = ((NEW - OLD) / OLD) * 100
% = ((0.417 - 0.293) / 0.293 = 42.32

*IF YOU HAVE ANY PROBLEMS, JUST LET ME KNOW IN THE COMMENTS AND I WILL TRY AND HELP